A researcher estimated the following equations in the demand and supply of shirts. D=200-25p S=16+11p. Calculate the equilibrium price and quantity.

1
Expert's answer
2022-03-22T11:08:05-0400

at equilibrium D=S,where

D=200-25p and S=16+11p

therefore:

200-25p=16+11p

200-16=25p+11p

184=36p

p=5.11


Dq=200-25p

=200-(25"\times5.11)"

=200-127.75

Q=72.25

at equilibrium ,price=5.11,quantity=72.25